OpenStudy (anonymous):

Please help! If 3x^2-2x+7=0 then (x-1/3)^2

OpenStudy (kropot72):

First you need to solve the quadratic and then evaluate the expression \[\large (x-\frac{1}{3})^{2}\] Is that the case?

OpenStudy (anonymous):

how do I solve it>? @kropot72

OpenStudy (kropot72):

You need to use the formula that will solve any quadratic: \[\large x=\frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\] I this case: a = 3; b = -2 and c = 7.

OpenStudy (anonymous):

ok Thanks :)

OpenStudy (kropot72):

Note that there is no real solution to the quadratic, the reason being that the discriminant is the square root of -80. Therefore the solution must be imaginary.

OpenStudy (anonymous):

what do you mean with imaginary

OpenStudy (anonymous):

I got -80?

OpenStudy (kropot72):

The square root of a negative number cannot be a real number. Are you sure that you typed the quadratic correctly? Yes the value of the discriminant is -80. When you take the square root, this is the result: \[\large \sqrt{-80}=\sqrt{-1}\sqrt{80}=i \sqrt{80}\] where i represents the square root of negative 1.

OpenStudy (anonymous):

yeah

OpenStudy (kropot72):

So the solutions to the quadratic are: \[\large x=\frac{2+8.94i}{6},\ x=\frac{2-8.94i}{6}\]

OpenStudy (kropot72):

These solutions can be simplified by dividing each term in the numerator by 6.

OpenStudy (kropot72):

If you have not studied imaginary numbers, perhaps this question is not appropriate for you at this time.

OpenStudy (kropot72):

Do you have answer choices?

OpenStudy (anonymous):

yeah its A. 20/9 B. 7/9 C. -7/9 D. -8/9 E. -20/9

OpenStudy (kropot72):

One of these solutions is correct, which confirms my understanding of the question.

OpenStudy (anonymous):

which is?

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

@kropot72 I'm confused on this ._.

OpenStudy (kropot72):

To get an exact solution it is best to put the square root of 80 in this form: \[\large \sqrt{80}=4\sqrt{5}\] Then the solutions of the quadratic become: \[\large x=\frac{1}{3}+\frac{2\sqrt{5}i}{3},\ \frac{1}{3}-\frac{2\sqrt{5}i}{3}\] When each of these two values of x are plugged into the expression \[\large (x-\frac{1}{3})^{2}\] we get: \[\large (\frac{1}{3}+\frac{2\sqrt{5}i}{3}-\frac{1}{3})^{2}=(\frac{2\sqrt{5}i}{3})^{2}=-\frac{20}{9}\] and \[\large (\frac{1}{3}-\frac{2\sqrt{5}i}{3}-\frac{1}{3})=(-\frac{2\sqrt{5}i}{3})^{2}=-\frac{20}{9}\]